for what values of p is 4x^2+px+49=0 a perfect square?
Answer by checkley77(12844)

4x^2+px+49=0
p=+/-28
(4x^2+/-28x+49)=0
Proof:
2x+7
2x+7
---------------
4x^2+14x+14x+49=4x^2+28x+49
2x-7
2x-7
------------------
4x^2-14x-14x+49=4x^2-28x+49